# CRAFT-LoRA: Content-Style Personalization via Rank-Constrained Adaptation and Training-Free Fusion

### v1

@@ 4.1 Implementation Details
For content concepts, we use the DreamBooth dataset  [27], containing 30 themes with 4-5 images each. Style concepts come from the StyleDrop  [33] image set, supplemented with classical and contemporary art style images. Importantly, we train each LoRA module (content or style) using only one reference image. To enhance content-style separation during backbone fine-tuning, we construct specialized datasets from our main content and style sources. These datasets emphasize either content preservation or style transfer properties through a controlled generation process based on frequency separation, as content typically resides in lower frequencies and style in higher frequencies. This approach creates clear content-style pairs that provide disentangled signals during backbone fine-tuning, with details provided in the Appendix.Appendix.
All experiments use the Stable Diffusion XL base model  [25] on NVIDIA 4090 GPUs. Our evaluation includes both locally trained LoRAs following our procedure and publicly available pre-trained LoRAs.For locally trained LoRAs, we implement DreamBooth training with rank ($r$) = 64, optimizing weights using Adam-8bit for 1000 steps with batch size 1 and learning rate $1\text{\times}{10}^{-5}$. Training uses minimal prompts like ”A [identifier] photo”, where ”[identifier]” (i.e. ”[c]” for content, ”[s]” for style) serves as a placeholder token representing the specific concept.
@@
(b) Ablation study over key components.
Rank-limited Backbone Fine-tuning
Setting Content Sim. Style Sim. Cross-Influence Score $\downarrow$
w/o 0.73 0.69 0.172
w/ 0.82 0.78 0.087
Asymmetric Classifier-Free Guidance
Setting Content Sim. Style Sim. Combination Score $\uparrow$
w/o 0.77 0.64 0.75
w/ 0.84 0.78 0.88
@@
Direct Merging 45.0% 40.0% 42.5%
StyleDrop 55.0% 72.5% 60.0%
ZipLoRA 62.5% 65.0% 67.5%
KLoRA 65.0% 70.0% 70.0%
BLoRA 70.0% 67.5% 72.5%
Ours 82.5% 85.0% 87.5%
Figure 6: Quantitative evaluation of our method. (a) Comparison against baselines using CLIP-I similarities and GPT-4o scores. (b) Ablation study analyzing the effects of Rank-limited Backbone Fine-tuning and Asymmetric CFG. (c) User study results on a 1–5 scale (30 participants, 50 samples), showing clear preference for our method across content fidelity, style fidelity, and overall coherence.
@@ 4.2 Results and Analysis
Visual Results. Figure 4 compares our method with representative LoRA fusion baselines, including ZipLoRA, BLoRA, and KLoRA. Prompts are formatted as “A [content] <c> in [style] <s>” with different content–style pairs such as dogs, plush toys, and objects. While existing methods often compromise either content fidelity e.g. structural distortion or style rendering e.g. muted patterns, our approach produces outputs that preserve object identity while faithfully capturing artistic styles. Figure 5 provides extended qualitative demonstrations. The left part shows content–style combinations augmented with additional prompt descriptions e.g. “catching a frisbee”, “playing with a ball”, or “driving a car”. Our method integrates these dynamic semantics while faithfully maintaining the designated content and style. The right part isolates the effect of single-branch adapters. Activating only the content branch preserves structural fidelity of objects across neutral renderings, while activating only the style branch applies stylistic transformations consistently across varied subjects. These observations confirm that our framework achieves effective disentanglement and controllability during inference.
@@
Quantitative Evaluation. We evaluate our method using both automatic and human-centered metrics. For automatic evaluation, we adopt Content Similarity and Style Similarity based on CLIP-I embeddings, as well as a Combination Score derived from GPT-4o that measures the overall coherence of content–style integration. Each score is averaged over 10 distinct content–style pairs, with 10 generated images per pair. As shown in Table 6(a), baseline methods exhibit different trade-offs: Direct Merging fails to preserve either content or style, StyleDrop emphasizes style at the expense of structure, and ZipLoRA, KLoRA, and BLoRA achieve more balanced results but remain limited in overall coherence. Our method consistently achieves the highest scores across all metrics, demonstrating its ability to preserve structural fidelity while maintaining faithful style transfer and coherent integration.
Ablation and User Study. Table 6(b) analyzes the contributions of our two key components. Rank-limited backbone fine-tuning improves CLIP-I content/style similarity and reduces cross-influence, indicating stronger content–style disentanglement. The Cross-Influence score measures the relative drop in content/style similarity when composing both factors versus single-factor baselines. Specifically, it is computed as the product of CLIP-I similarities between the generated image and both reference images; lower values indicate better disentanglement with reduced mutual interference. Asymmetric CFG further strengthens integration quality, reflected by higher human-aligned combination scores and consistent gains in CLIP-I. Table 6(c) further reports results from a user study with 30 participants. Participants consistently preferred our method across content fidelity, style fidelity, and overall coherence, aligning with the automatic metrics. This combination of quantitative and perceptual evidence confirms that our framework achieves superior disentanglement, fidelity, and integration compared with prior approaches.
@@ 5 Conclusion
We presented a framework for personalized image generation that integrates efficient adaptation strategies. Rank-constrained backbone fine-tuning encourages distinct subspaces, providing a inductive bias for disentanglement. An expert encoder with selective aggregation enhances semantic guidance, enabling flexible composition, and timestep-dependent asymmetric CFG introduces a training-free sampling strategy that improves visual quality.
Our experiments substantiate these contributions, demonstrating a measurable reduction in content–style cross-influence, consistent improvements across automatic evaluation metrics, and clear user preference in perceptual studies. Limitations include architecture-dependent layer selection, reliance on text embedding quality for semantic extension, and minor overhead from timestep-dependent correction. Future work will explore automated layer assignment and extensions to broader diffusion models and modalities.

### v2

@@ 4.1 Implementation Details
For content concepts, we use the DreamBooth dataset [30] (30 themes, 4-5 images each). Style concepts come from StyleDrop [37], supplemented with art style images. Each LoRA module is trained from a single reference image. We construct 100 contrastive pairs from these sources via frequency-domain decomposition to enhance content-style separation during backbone fine-tuning. For evaluation, we use strictly disjoint train/test splits. Each metric is averaged over 10 content-style pairs with 10 generated images per pair. Content and Style Similarity are computed using CLIP-I embeddings. The Combination Score follows MLLM-based binary judgment practices [19]: GPT-4o is presented with the content reference, style reference, and output, and asked whether the output successfully integrates both content identity and artistic style. The final score is the average success rate over 50 evaluations per sample.
[…]
@@
(b) Component ablation study
Rank-FT Router ACFG Content Sim. $\uparrow$ Style Sim. $\uparrow$ Comb. Score $\uparrow$
0.65 0.60 0.62
0.73 0.70 0.72
0.69 0.65 0.67
0.68 0.64 0.66
0.76 0.76 0.80
0.75 0.73 0.78
0.71 0.68 0.70
0.79 0.80 0.83
@@
Direct Merging 2.3 2.0 2.1
StyleDrop 2.8 3.6 3.0
ZipLoRA 3.1 3.3 3.4
KLoRA 3.3 3.5 3.5
BLoRA 3.5 3.4 3.6
Ours 4.1 4.3 4.4
Figure 6: Quantitative evaluation of our method. (a) Comparison against baselines using CLIP-I similarities and GPT-4o binary judgment scores. (b) Component ablation: Rank-FT contributes most to disentanglement, Router provides content-style routing, and ACFG stabilizes fusion; all three are complementary. The baseline (no components) corresponds to Direct Merging. (c) User study ratings on a 1–5 scale (30 participants, 50 samples).
@@ 4.2 Results and Analysis
Visual Results. Figure 4 compares our method with representative LoRA fusion baselines, including ZipLoRA, BLoRA, and KLoRA. Prompts are formatted as “A [content] <c> in [style] <s>” with different content–style pairs such as dogs, plush toys, and objects. While existing methods often compromise either content fidelity e.g. structural distortion or style rendering e.g. muted patterns, our approach produces outputs that preserve object identity while faithfully capturing artistic styles. Figure 5 provides extended qualitative demonstrations. The top part shows content–style combinations augmented with additional prompt descriptions e.g. “catching a frisbee”, “playing with a ball”, or “driving a car”. Our method integrates these dynamic semantics while faithfully maintaining the designated content and style. The bottom part isolates the effect of single-branch adapters. Activating only the content branch preserves structural fidelity of objects across neutral renderings, while activating only the style branch applies stylistic transformations consistently across varied subjects. These observations confirm that our framework achieves effective disentanglement and controllability during inference.
@@
Quantitative Evaluation. We evaluate our method using both automatic and human-centered metrics. For automatic evaluation, we adopt Content Similarity and Style Similarity based on CLIP-I embeddings, as well as a Combination Score derived from GPT-4o that measures the overall coherence of content–style integration. Each score is averaged over 10 distinct content–style pairs, with 10 generated images per pair. As shown in Table 6 (a), baseline methods exhibit different trade-offs: Direct Merging fails to preserve either content or style, StyleDrop emphasizes style at the expense of structure, and ZipLoRA, KLoRA, BLoRA, and LoRA.rar achieve more balanced results but remain limited in overall coherence. Our method consistently achieves the highest scores across all metrics, demonstrating its ability to preserve structural fidelity while maintaining faithful style transfer and coherent integration.
Ablation and User Study. Table 6 (b) presents a comprehensive component ablation. Among individual components, Rank-FT contributes most significantly, improving Content Sim. by +0.08 and Style Sim. by +0.10 over the baseline, confirming its central role in content-style disentanglement. The Router provides complementary gains through layer-wise content-style routing, while ACFG stabilizes fusion quality. Combining any two components yields further improvements, and the full configuration achieves the best performance across all metrics, demonstrating that the three components are complementary.
Table 6 (c) reports results from a user study with 30 participants rating 50 samples on a 1–5 scale. Our method achieves the highest ratings across content fidelity, style fidelity, and overall coherence, consistent with the automatic metrics. This combination of quantitative and perceptual evidence confirms that our framework achieves superior disentanglement, fidelity, and integration compared with prior approaches.
Failure Cases Analysis. Despite the effectiveness of our approach, we identified three main failure modes during the diverse implementation process. First, when the reference image contains highly entangled content and style (e.g. a cartoon character whose identity is inseparable from the rendering), the method struggles to isolate content from style, resulting in style leakage into the content branch. Second, for styles dominated by low-frequency characteristics such as flat color palettes, the frequency-based separation misassigns these features to the content domain, leading to incomplete style capture. Third, abstract style descriptions (e.g. “ethereal mood”) may lack precise CLIP embeddings, resulting in weak or inconsistent style transfer. Adjusting the frequency cutoff $\sigma$ or adopting stronger encoders could mitigate these issues.
@@ 5 Conclusion
We presented a framework for personalized image generation that integrates efficient adaptation strategies. Rank-constrained backbone fine-tuning encourages distinct subspaces, providing an inductive bias for disentanglement. An expert encoder with selective aggregation enhances semantic guidance, enabling flexible composition, and timestep-dependent asymmetric CFG introduces a training-free sampling strategy that improves visual quality. Our experiments demonstrate measurable reduction in content-style cross-influence, consistent improvements across automatic metrics, and clear user preference in perceptual studies. Limitations include architecture-dependent layer selection, reliance on text embedding quality, frequency-based separation assumptions that may not hold for flat textures, and the two-branch ACFG structure which limits multi-content or multi-style mixing. Standard SDXL LoRAs can be used with ACFG directly on the base model, though full disentanglement benefits require retraining on the fine-tuned backbone. Future work will explore automated layer assignment, multi-concept scheduling, and extensions to architectures such as KOALA [16] and SANA [42].